OpenStudy (mathmath333):

Aman and Biman started a walkathon around a circular track starting from the same point on the track in the opposite directions.They meet for the first time after time '\(t\)'.Had they walked in the same direction with their speeds intact ,they would have met after time '\(7t\)'.It was also observed that Aman was slower that Biman. Aman's speed was measured at \(12\) kmph.find the speed of Biman. a.) \(4\) m/s. b.) \(5\) m/s. c.) \(6\) m/s. d.) None of these.

OpenStudy (unklerhaukus):

if they have met when they are traveling the same direction, one of them must have completed exactly one more lap than the other

OpenStudy (mathmath333):

yes.

OpenStudy (mathmath333):

\(\large \color{black}{ t=\dfrac{d}{x+12}\\\\\\ 7t=\dfrac{d}{x-12}\\\\\\\ ? \hspace{.33em}\\~\\ }\)

OpenStudy (mathmath333):

let \(\large \color{black}{x \hspace{.33em}\\~\\ }\) be the speed of Biman

OpenStudy (mathmath333):

ohk got the answer.

OpenStudy (mathmath333):

@rational what did u get.

OpenStudy (rational):

hey why do we even need to do any work.. look at the options, none of them are greater than 12.. so the answer would be simply "None of these"

OpenStudy (unklerhaukus):

but the units/

OpenStudy (rational):

Oops!

OpenStudy (unklerhaukus):

12 [km/hr] = 3.3 [m/s]

OpenStudy (rational):

im getting 16kmph for Biman

OpenStudy (mathmath333):

16 is much bigger than options

OpenStudy (mathmath333):

but it is correct

OpenStudy (rational):

16kmph = 4.44 m/s

OpenStudy (mathmath333):

anyways whether u change unit or not

OpenStudy (rational):

btw i have just divided your both equations to eliminate t and d http://gyazo.com/5a53d8038fccf685069e0b21fb8ef084

OpenStudy (mathmath333):

at first i forgot about units , and was worried about the bigger value of \(16\)

OpenStudy (rational):

haha so the key thing to notice when running in same direction is `if they have met when they are traveling the same direction, one of them must have completed exactly one more lap than the other`